# Deep Learning the Forecast of Galactic Cosmic-Rays Spectra

### v2

@@ Abstract
Acknowledgments.—We appreciate the insightful discussions with Weiwei Xu, Ran Huo, Kai Zhou, Zhaoming Wang, Yao Chen, and Siqi Wang. This work is supported by the Taishan Scholars Program of Shandong Province under Grant No. tsqnz20221162 (Y. D.) and Grant No. 202103143 (X. L.), the Shandong Excellent Young Scientists Fund Program (Overseas) under Grant No. 2023HWYQ-106 (Y. D.), the NSFC under Grant No. U2106201 (X. L.) and Grant No. 42404172 (X. S.), and the Natural Science Foundation of Shandong Province under Grant No. ZR2022QD033 (X. S.).
Appendix A Supplemental Material
Figure 4 illustrates daily solar activity data, spanning observed values from 2011 to 2019 and predicted values from 2019 to 2031, which are used as inputs to the LSTM network. Figures 5 and 6 display the training, testing, and prediction results for Helium flux in the [2.97, 3.29] GV rigidity range one day ahead. These figures compare models trained with five solar parameters alone and those augmented with the historical Helium flux as an additional input. Finally, Figure 7 highlights the LSTM neural network’s averaged near-future prediction on Helium flux across different model configurations.
Figure 4: Daily solar activity observed from 2011 to 2019 and predicted from 2019 to 2031 are shown as inputs to the LSTM neural network, plotted over time. The parameters displayed include the heliospheric magnetic field (HMF) ($B$), solar wind speed ($v_{sw}$) near Earth, heliospheric current sheet (HCS) tilt angle ($\alpha$), solar polarity (A), and sunspot numbers (SSN).

### v3

@@ 4 Conclusions
Acknowledgments
We appreciate the insightful discussions with Weiwei Xu, Ran Huo, Kai Zhou, Zhaoming Wang, Yao Chen, and Siqi Wang. This work is supported by the Taishan Scholars Program of Shandong Province under Grant No. tsqnz20221162 (Y. D.) and Grant No. 202103143 (X. L.), the Shandong Excellent Young Scientists Fund Program (Overseas) under Grant No. 2023HWYQ-106 (Y. D.), the NSFC under Grant No. U2106201 (X. L.) and Grant No. 42404172 (X. S.), and the Natural Science Foundation of Shandong Province under Grant No. ZR2022QD033 (X. S.).
appendix
@@ Acknowledgments
Figures. 6 and 7 display the training, testing, and prediction results for Helium flux in the [2.97, 3.29] GV rigidity range one day ahead. These figures compare models trained with five solar parameters alone and those augmented with the historical Helium flux as an additional input.